ifundefineddocumentclass@Author@uthor@true@german@erman@true@ancient@layout@ncient@layout@false@file ¸@enumi ¸@enumi(lph ¸@enumii) ¸@enumiii ¸@tocdepth ¿-2empdima3em@ numwidth numwidth 1@xt@numwidth

2 obreaktrue ¸@tocdepth ¿@ne

lus@@ empdima1.5em@ numwidth numwidth pdima 1  @xt@numwidth  2

ighpenalty